OpenStudy (anonymous):

Least To Greatest 1/2 , 3/7 , 7/10

OpenStudy (anonymous):

3/7, 1/2, 7/10

OpenStudy (anonymous):

I meant to put 3/8

OpenStudy (anonymous):

either way it is teh least

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[\left\{\frac{3}{8},\frac{3}{7} ,\frac{1}{2} ,\frac{7}{10}\right\}=\{0.375,0.428571,0.5,0.7\} \]